# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 9.1 Robust Control of Systems Modeled by DNNs
to cancel out the unknown term $r(x,u,t)$ of the dynamical system (9.7) by the model $r_{L}(x,u,t)$ learned using trajectory data, where $u^{*}$ is a nominal stabilizing control input for $\dot{x}=f(x,t)+B(x,t)u$ given by, e.g., Theorems 4.2 and 4.6. However, the equation (9.8) depends implicitly on $u$, which brings extra computational burden especially if the learned model $r_{L}(x,u,t)$ is highly nonlinear as in DNNs. In Shi et al. [2019], a discrete-time nonlinear controller is proposed to iteratively solve (9.8).
@@
where $\Delta u=u-u^{\prime}$. Thus, the assumption $L_{u}<1$ ensures that $\mathcal{F}$ is a contraction mapping for fixed $x,t$ Shi et al. [2019]. $\qed$
By applying contraction theory to the discrete-time controller (9.9) of Lemma 9.1, we can guarantee the stability of (9.7) if $r_{L}$ is modeled by a spectrally-normalized DNN of Definition 6.3.
@@
Let $x$ be the trajectory of the following ideal system without the unknown part $r$ of the dynamics (9.7):
@@
where $u^{*}$ is a stabilizing controller that renders (9.11) contracting as in Theorem 9.1 for $M$ which satisfies $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32). Note that such $u^{*}$ can be designed by using, e.g., Theorems 4.2 and 4.6. Suppose that the true dynamics (9.7) is controlled by (9.9) and
@@
for $x^{*}$ in (9.7Shi et al. [2019]. If $\exists\bar{b}\in\mathbb{R}_{\geq 0}$ s.t. $\|B(x,t)\|\leq\bar{b}$, and if $r_{L}$ is modeled by a spectrally-normalized DNN of Definition 6.3 to have
@@
where $V_{\ell}=\int^{x}_{x^{*}}\|\Theta\delta x\|$ as in Theorem 2.3 with $M=\Theta^{\top}\Theta$. Furthermore, the system (9.7) with the controller (9.9) is robust against deterministic and stochastic disturbances.
@@
where (9.12) is used to obtain the second inequality. Since we have $\|B(x,t)\|\leq\bar{b}$ and the closed-loop system $\dot{x}=f_{cl}(x,t)$ is contracting, Theorem 5.2 holds with $\epsilon_{\ell 0}=\bar{b}\epsilon_{\ell}$ and $\epsilon_{\ell 1}=\bar{b}L_{u}\rho$ as long as we select $L_{u}$ to satisfy (9.15), resulting in the bound (9.14). The final robustness statement follows from Theorems 5.2 and 5.3. $\qed$
Theorem 9.2 implies that the control synthesis algorithms via contraction theory, including robust control of Theorems 4.2 and 4.6 (CV-STEM), learning-based robust control of Theorems 6.1, 6.36.5, and 7.1 (NCM, LAG-ROS), can be enhanced to provide explicit robustness and stability guarantees even for systems partially modeled by DNNs that depend nonlinearly on $u$.
@@
Let us consider the following Lagrangian system of Example 2.6 perturbed externally by unknown control non-affine residual disturbance:
@@
where $x=[\mathtt{q}^{\top},\dot{\mathtt{q}}^{\top}]^{\top}$ and the other variables are as given in Example 2.6. Using the result of Theorem 9.2, we can design a discrete-time nonlinear controller by augmenting the exponentially stabilizing controller of Example 2.6 with a learned residual part $r_{L}(x,u)$ as follows:
@@
where $\dot{\mathtt{q}}_{r}=\dot{\mathtt{q}}_{d}(t)-\Lambda(t)(\mathtt{q}-\mathtt{q}_{d}(t))$, $\mathcal{K}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times n}$, $\Lambda:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times n}$, and $(\mathtt{q}_{d},\dot{\mathtt{q}}_{d})$ is the target trajectory of the state $(\mathtt{q},\dot{\mathtt{q}})$, and $\mathcal{K},\Lambda\succ 0$ are control gain matrices (design parameters). Again, note that $\dot{\mathcal{H}}-2\mathcal{C}$ is skew-symmetric with $\mathcal{H}\succ 0$ by construction. This gives us the following closed-loop virtual system of a smooth path $q(\mu,t)$ parameterized by $\mu\in[0,1]$, which has $q(\mu=0,t)=\dot{\mathtt{q}}_{r}$ and $q(\mu=1,t)=\dot{\mathtt{q}}$ as its particular solutions as in Example 2.6, but now with non-zero perturbation due to $r(x,u)$:
@@
After some algebra as in the proof of Theorem 9.2, we can show that
@@
where $\mathcal{H}=\Theta^{\top}\Theta$, $h_{\ell}\mathrm{I}\preceq\mathcal{H}\preceq h_{u}\mathrm{I}$, $k_{\ell}\mathrm{I}\preceq\mathcal{K}$, $\|u_{k}-u_{k-1}\|\leq\rho\|\dot{\mathtt{q}}-\dot{\mathtt{q}}_{r}\|$, $\epsilon_{\ell}$ is the learning error of $r_{L}$, and $L_{u}$ is the Lipschitz constant of $r_{L}$ (assuming $r_{L}$ is modeled by a spectrally-normalized DNN of Definition 6.3). This indeed indicates that the tracking error of the Lagrangian system (9.18) is exponentially bounded as proven in Theorem 9.2.
In Shi et al. [2019], the technique in Theorem 9.2 and in Example 9.1 is used to perform precise near-ground trajectory control of multi-rotor drones, by learning complex aerodynamic effects caused by high-order interactions between multi-rotor airflow and the environment. It is demonstrated that it significantly outperforms a baseline nonlinear tracking controller in both landing and cross-table trajectory tracking tasks. Theorem 9.2 enables applying it to general nonlinear systems with state and control dependent uncertainty, as long as we have a nominal exponentially stabilizing controller (which can be designed by Theorems 4.2 and 4.6, or approximately by Theorems 6.1, 6.36.5, and 7.1).
9.2 Learning Contraction Metrics from Trajectory Data
@@
Table 6: Notations in Sec. 9.2 Boffi et al. [2020].
@@
The following theorem states that a contraction metric defined by $\mathcal{M}$, learned from trajectory data, indeed guarantees contraction of the unknown underlying dynamical system Boffi et al. [2020].
@@
Suppose that $\mathcal{S}_{s}\subseteq\mathbb{R}^{n}$ is full-dimensional, $n\geq 2$, and $\dot{x}=f(x)$ is contracting in a metric defined by a uniformly bounded $M(x)\succ 0$, i.e.,
@@
See Theorem 5.2 of Boffi et al. [2020]. $\qed$
Agreeing with our intuition, Theorem 9.3 ensures that the region $\tilde{S}(r_{b})$ where the contraction condition holds becomes larger, as the upper bound $\epsilon_{\ell}$ on $\nu(Z_{b})$ for the contraction violation set $Z_{b}$ becomes smaller, yielding the learned contraction metric of higher quality. In Boffi et al. [2020], it is proven that if the samples are drawn uniformly from $\mathcal{S}_{s}\times\mathbb{S}^{n-1}$, we have $\epsilon_{\ell}\leq O(k\cdot\text{polylog}(N)/N)$ decay rates for various parametric and nonparametric function classes, where $N$ is the number of sampled trajectories and $k$ is the effective number of parameters of the function class of interest. Such a learned contraction metric can be used in Theorems 9.1 and 9.2 for certifying robustness and stability of model-free systems.
10 Concluding Remarks

### v8

@@ 9.1. Robust Control of Systems Modeled by DNNs
to cancel out the unknown term $r(x,u,t)$ of the dynamical system (9.7) by the model $r_{L}(x,u,t)$ learned using trajectory data, where $u^{*}$ is a nominal stabilizing control input for $\dot{x}=f(x,t)+B(x,t)u$ given by, e.g., Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:ccm_cvstem. However, the equation (9.8) depends implicitly on $u$, which brings extra computational burden especially if the learned model $r_{L}(x,u,t)$ is highly nonlinear as in DNNs. In neurallander, a discrete-time nonlinear controller is proposed to iteratively solve (9.8).
@@
where $\Delta u=u-u^{\prime}$. Thus, the assumption $L_{u}<1$ ensures that $\mathcal{F}$ is a contraction mapping for fixed $x,t$ neurallander. $\qed$
By applying contraction theory to the discrete-time controller (9.9) of Lemma 9.1, we can guarantee the stability of (9.7) if $r_{L}$ is modeled by a spectrally-normalized DNN of Definition LABEL:DEF:Def:SN.
@@
Robust Nonlinear Control of Learned ModelsThm:neurallander Let $x$ be the trajectory of the following ideal system without the unknown part $r$ of the dynamics (9.7):
@@
where $u^{*}$ is a stabilizing controller that renders (9.11) contracting as in Theorem LABEL:THM:Thm:NCMstability_modelfree for $M$ which satisfies $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32). Note that such $u^{*}$ can be designed by using, e.g., Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:ccm_cvstem. Suppose that the true dynamics (9.7) is controlled by (9.9) and
@@
for $x^{*}$ in (9.7neurallander. If $\exists\bar{b}\in\mathbb{R}_{\geq 0}$ s.t. $\|B(x,t)\|\leq\bar{b}$, and if $r_{L}$ is modeled by a spectrally-normalized DNN of Definition LABEL:DEF:Def:SN to have
@@
where $V_{\ell}=\int^{x}_{x^{*}}\|\Theta\delta x\|$ as in Theorem LABEL:THM:Thm:path_integral with $M=\Theta^{\top}\Theta$. Furthermore, the system (9.7) with the controller (9.9) is robust against deterministic and stochastic disturbances.
@@
where (9.12) is used to obtain the second inequality. Since we have $\|B(x,t)\|\leq\bar{b}$ and the closed-loop system $\dot{x}=f_{cl}(x,t)$ is contracting, Theorem LABEL:THM:Thm:contraction_learning holds with $\epsilon_{\ell 0}=\bar{b}\epsilon_{\ell}$ and $\epsilon_{\ell 1}=\bar{b}L_{u}\rho$ as long as we select $L_{u}$ to satisfy (9.15), resulting in the bound (9.14). The final robustness statement follows from Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto. $\qed$
Theorem LABEL:THM:Thm:neurallander implies that the control synthesis algorithms via contraction theory, including robust control of Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:ccm_cvstem (CV-STEM), learning-based robust control of Theorems LABEL:THM:Thm:NCMstability1, LABEL:THM:Thm:NCMstability2LABEL:THM:Thm:ncm_ccm_clf, and LABEL:THM:Thm:lagros_stability (NCM, LAG-ROS), can be enhanced to provide explicit robustness and stability guarantees even for systems partially modeled by DNNs that depend nonlinearly on $u$.
@@
Lagrangian Systems with Learned Residualsex:neural_lander Let us consider the following Lagrangian system of Example LABEL:EX:ex:lag_metric perturbed externally by unknown control non-affine residual disturbance:
@@
where $x=[\mathtt{q}^{\top},\dot{\mathtt{q}}^{\top}]^{\top}$ and the other variables are as given in Example LABEL:EX:ex:lag_metric. Using the result of Theorem LABEL:THM:Thm:neurallander, we can design a discrete-time nonlinear controller by augmenting the exponentially stabilizing controller of Example LABEL:EX:ex:lag_metric with a learned residual part $r_{L}(x,u)$ as follows:
@@
where $\dot{\mathtt{q}}_{r}=\dot{\mathtt{q}}_{d}(t)-\Lambda(t)(\mathtt{q}-\mathtt{q}_{d}(t))$, $\mathcal{K}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times n}$, $\Lambda:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times n}$, and $(\mathtt{q}_{d},\dot{\mathtt{q}}_{d})$ is the target trajectory of the state $(\mathtt{q},\dot{\mathtt{q}})$, and $\mathcal{K},\Lambda\succ 0$ are control gain matrices (design parameters). Again, note that $\dot{\mathcal{H}}-2\mathcal{C}$ is skew-symmetric with $\mathcal{H}\succ 0$ by construction. This gives us the following closed-loop virtual system of a smooth path $q(\mu,t)$ parameterized by $\mu\in[0,1]$, which has $q(\mu=0,t)=\dot{\mathtt{q}}_{r}$ and $q(\mu=1,t)=\dot{\mathtt{q}}$ as its particular solutions as in Example LABEL:EX:ex:lag_metric, but now with non-zero perturbation due to $r(x,u)$:
@@
After some algebra as in the proof of Theorem LABEL:THM:Thm:neurallander, we can show that
@@
where $\mathcal{H}=\Theta^{\top}\Theta$, $h_{\ell}\mathrm{I}\preceq\mathcal{H}\preceq h_{u}\mathrm{I}$, $k_{\ell}\mathrm{I}\preceq\mathcal{K}$, $\|u_{k}-u_{k-1}\|\leq\rho\|\dot{\mathtt{q}}-\dot{\mathtt{q}}_{r}\|$, $\epsilon_{\ell}$ is the learning error of $r_{L}$, and $L_{u}$ is the Lipschitz constant of $r_{L}$ (assuming $r_{L}$ is modeled by a spectrally-normalized DNN of Definition LABEL:DEF:Def:SN). This indeed indicates that the tracking error of the Lagrangian system (9.18) is exponentially bounded as proven in Theorem LABEL:THM:Thm:neurallander.
In neurallander, the technique in Theorem LABEL:THM:Thm:neurallander and in Example LABEL:EX:ex:neural_lander is used to perform precise near-ground trajectory control of multi-rotor drones, by learning complex aerodynamic effects caused by high-order interactions between multi-rotor airflow and the environment. It is demonstrated that it significantly outperforms a baseline nonlinear tracking controller in both landing and cross-table trajectory tracking tasks. Theorem LABEL:THM:Thm:neurallander enables applying it to general nonlinear systems with state and control dependent uncertainty, as long as we have a nominal exponentially stabilizing controller (which can be designed by Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:ccm_cvstem, or approximately by Theorems LABEL:THM:Thm:NCMstability1, LABEL:THM:Thm:NCMstability2LABEL:THM:Thm:ncm_ccm_clf, and LABEL:THM:Thm:lagros_stability).
9.2. Learning Contraction Metrics from Trajectory Data
@@
Table 6: Notations in Sec. 9.2 boffi2020learning.
@@
The following theorem states that a contraction metric defined by $\mathcal{M}$, learned from trajectory data, indeed guarantees contraction of the unknown underlying dynamical system boffi2020learning.
@@
Contraction in Learned Contraction MetricsThm:data_driven_cont_thm Suppose that $\mathcal{S}_{s}\subseteq\mathbb{R}^{n}$ is full-dimensional, $n\geq 2$, and $\dot{x}=f(x)$ is contracting in a metric defined by a uniformly bounded $M(x)\succ 0$, i.e.,
@@
See Theorem 5.2 of boffi2020learning. $\qed$
Agreeing with our intuition, Theorem LABEL:THM:Thm:data_driven_cont_thm ensures that the region $\tilde{S}(r_{b})$ where the contraction condition holds becomes larger, as the upper bound $\epsilon_{\ell}$ on $\nu(Z_{b})$ for the contraction violation set $Z_{b}$ becomes smaller, yielding the learned contraction metric of higher quality. In boffi2020learning, it is proven that if the samples are drawn uniformly from $\mathcal{S}_{s}\times\mathbb{S}^{n-1}$, we have $\epsilon_{\ell}\leq O(k\cdot\text{polylog}(N)/N)$ decay rates for various parametric and nonparametric function classes, where $N$ is the number of sampled trajectories and $k$ is the effective number of parameters of the function class of interest. Such a learned contraction metric can be used in Theorems LABEL:THM:Thm:NCMstability_modelfree and LABEL:THM:Thm:neurallander for certifying robustness and stability of model-free systems.
10. Concluding Remarks